### v1

@@ 5 PDM systems invariant with respect to shifts
Let Q be an integral of motion (7) admitted by equation (1) with arbitrary elements (9). By definition $P_{3}$ is the integral of motion too, the same is true for the commutators $[P_{3},Q],[P_{3},[P_{3},Q]$ and $[P_{3},[P_{3},[P_{3},Q]].$ Thus any second order symmetry induces the symmetry generated by $\mu^{ab}$ independent on $x_{3}$ which are itemized below together with one additional special tensor:
@@
Classifying PDM systems admitting these induced symmetries we can obtain the complete list of the systems invariant with respect to shifts and admitting second order integrals of motion. Then it will be necessary to specify the additional integrals of motion which can be admitted by the found systems. As a result we obtain the complete list of the systems presented in the following Tables 3-5. The detailed calculations are presented in Sections 5.1- 5.4.
5.1 Symmetries whose differential part $Q_{(n)}$ (36) does not depend on $x_{1},x_{2}$ and $x_{3}$
The first step of our analysis is to find the systems with the special property specified in the subsection title. The corresponding Killing tensors are specified in the items presented in the above.
Let us start with the constant Killing tensors $\mu^{ab}=\mu^{ab}_{1}=\lambda_{1}^{ab}$. The related integrals of motion look as follows:
$Q_{1}=\lambda^{ab}P_{a}P_{b}+\left(N\cdot H\right)+\eta$ (85)
where we omit the subindex 1 for $\lambda_{1}^{ab}$.
Up to rotations around the third coordinate axis and the arbitrariness of function $g({\bf x})$ we can restrict ourselves to the following nonzero entries: $\lambda^{3\alpha},\alpha=1,2$, $\lambda^{11}=-\lambda^{22}=\mu$ and $\lambda^{12}=\lambda^{21}=\nu$ where $\mu$ and $\nu$ are arbitrary real numbers satisfying the condition $\mu\nu=0$. The corresponding operator (85) and determining equations (19) take the following forms:
$Q_{2}=\mu(P_{1}^{2}-P_{2}^{2})+\nu P_{1}P_{2}+\lambda^{31}P_{3}P_{1}+\lambda^{32}P_{3}P_{2}+\left(N\cdot H\right)+\eta$ (86)
and
$\begin{split}&\mu M_{1}+\nu M_{2}+N_{1}=0,\\ &\nu M_{1}-\mu M_{2}+N_{2}=0,\\ &\lambda^{31}M_{1}+\lambda^{32}M_{2}+N_{3}=0.\end{split}$ (87)
Let $\lambda^{31}=\lambda^{32}=\nu=0$ than solutions of equations (87) are
$M=F(x_{1})+G(x_{2}),\ N=\mu(-F(x_{1})+G(x_{2}))$ (88)
where $F(.)$ and $G(.)$ are arbitrary functions. The generic form of the related potential $V$ is the same as it has $M$ given in the above, but with another arbitrary functions, i.e.,
$V=\tilde{F}(x_{1})+\tilde{G}(x_{2})$ (89)
while the corresponding symmetry is reduced to the following form:
$Q_{2}=P_{1}^{2}-P_{2}^{2}+\left((-F(x_{1})+G(x_{2}))\cdot H\right)+\tilde{F}(x_{1})-\tilde{G}(x_{2}).$ (90)
Just this system is represented in Item 1 of Table 3.
In the case of trivial $\mu$ and $\nu$ and nontrivial $\lambda^{3\alpha}$ we can apply the rotations again and reduce $\lambda^{32}$ to zero. As a result we come to the following solution
$M=c_{1}x_{1}+G(x_{2}),\ V=c_{2}x_{1}+\tilde{G}(x_{2}),\ N=-\lambda^{31}c_{1}x_{3}.$ (91)
The corresponding symmetry (87) looks as:
$Q_{3}=2P_{3}P_{1}-(c_{1}x_{3}\cdot H)+c_{2}x_{3}.$ (92)
Notice that function $M$ in (91) is nothing but a particular case of the function presented in (88). It means that the PDM system with the mass and potential given in (91) admits both symmetries (89) and (92). This speculation is supported by the following observation: functions $M$ and $V$ fixed by equation (91) solve equations (87) where $\nu=\lambda^{33}=0$ but both $\mu$ and $\lambda^{31}$ are not trivial. The obtained solutions are presented in Item 2 of Table 4 where the additional symmetry of PDM hamiltonian with arbitrary elements (91) is represented also.
Finally, if all parameters $\mu,\lambda^{31}$ and $\lambda^{32}$ are nontrivial, the related solutions of (87) for the mass and potential are:
$M=c_{1}x_{1}+c_{2}x_{2},\ V=c_{3}x_{1}+c_{4}x_{2}+c_{5}$ (93)
and we have again a particular case of $M$ represented in (88) (and in (91)). In other words, the PDM Hamiltonian with the mass and potential given in (93) admits three independent integrals of motion which we represent in Item 2 of Table 4 together with the additional linearly independent but functionally dependent ones.
No $f$ $V$ Integrals of motion
2 $\frac{\tilde{r}^{2}}{F(\tilde{r})-G(\varphi)}$ $\frac{N(\varphi)+M(\tilde{r})}{G(\varphi)-F(\tilde{r})}$ $L_{3}^{2}+(G(\varphi)\cdot H)+N(\varphi)$
5 $\frac{\tilde{r}^{2}}{\lambda\tilde{r}^{2}+F(\varphi-\nu\ln(\tilde{r}))}$ $\frac{\kappa\tilde{r}^{2}+G(\varphi-\nu\ln(\tilde{r}))}{\lambda\tilde{r}^{2}+F(\varphi-\nu\ln(\tilde{r}))}$$\begin{array}[]{c}\{P_{3},D\}+\nu\{P_{3},L_{3}\}\\ -\left(2\lambda x_{3}\cdot H\right)+2\kappa x_{3}\end{array}$
6 $\frac{x_{2}^{2}}{x_{2}^{2}F(x_{1})+c_{1}}$ $\frac{x_{2}^{2}\tilde{F}(x_{1})+c_{2}}{x_{2}^{2}F(x_{1})+c_{1}}$ $\begin{array}[]{l}L_{1}^{2}-\left(\frac{c_{1}x_{3}^{2}}{x_{2}^{2}}\cdot H\right)+\frac{c_{2}x_{3}^{2}}{x_{2}^{2}},\\ \{P_{2},L_{1}\}+2\left(\frac{c_{1}x_{3}}{x_{2}^{2}}\cdot H\right)-2\frac{c_{2}x_{3}}{x_{2}^{2}},\\ P_{1}^{2}-P_{2}^{2}+\left((\frac{c_{1}}{x_{2}^{2}}-F(x_{1}))\cdot H\right)\\ -\frac{c_{2}}{x_{2}^{2}}+\tilde{F}(x_{1})\end{array}$
7 $\frac{\tilde{\omega}}{F(\tilde{\omega}+y)+G(\tilde{\omega}-y)}$ $\frac{\tilde{F}(\tilde{\omega}+y)+\tilde{G}(\tilde{\omega}-y)}{F(\tilde{\omega}+y)+G(\tilde{\omega}-y)}$ $\begin{array}[]{c}L_{3}^{2}\pm(P_{1}^{2}-P_{2}^{2})\\ +\left(g(F,G)\cdot H\right)-g(\tilde{F},\tilde{G})\end{array}$
8 $\begin{array}[]{c}\sum_{m}c_{m}\Omega_{m},\end{array}$ $\begin{array}[]{c}\frac{\sum_{m}\tilde{c}_{m}\Omega_{m}}{\sum_{m}c_{m}\Omega_{m}}\end{array}$ $\begin{array}[]{c}L_{3}^{2}\pm(P_{1}^{2}-P_{2}^{2})\\ -\left(y\sum_{m}c_{m}\Phi_{m-1}\cdot H\right)\\ +y\sum_{m}\tilde{c}_{m}\Omega_{m-1}\end{array}$
In Table 3 we use the following notations: $\Omega_{m}=\frac{(\tilde{\omega}+y)^{m}+(-1)^{m+1}(\tilde{\omega}-y)^{m}}{\tilde{\omega}},$ $y=x_{1}^{2}-x_{2}^{2}\mp 1,\ \tilde{\omega}=\sqrt{\tilde{r}^{4}-4y},c_{m}$ with $m=1,2,..$ are arbitrary parameters. This table represents all inequivalent PDM systems which admit two integrals of motion, one of which is the generator of shift transformations along the third coordinate axis and the other is a second order integral of motion. The two additional integrals of motion presented in Item 6 can be expressed via commutators of $P_{3}$ and $L_{1}^{2}$ and so are not independent. In other words the systems presented in Table 3 are integrable while superintegrable systems are represented in Table 4.
5.2 Integrals of motion whose differential part $Q_{(n)}$ (36) is independent on $x_{3}$ but depends on $x_{1}$ and $x_{2}$
@@ 5.1 Symmetries whose differential part $Q_{(n)}$ (36) does not depend on $x_{1},x_{2}$ and $x_{3}$
The second step of our programm is to classify the systems whose integrals of motion are generated by the Killing tensors specified at the beginning of Section 5, i.e., by the following linear combinations:
@@
$\begin{array}[]{l}Q=\omega(P_{1}^{2}-P_{2}^{2})+\kappa P_{1}P_{2}+\lambda^{3\alpha}P_{3}P_{\alpha}+\mu L_{3}P_{3}+\nu\{P_{1},L_{3}\}+\sigma L_{3}^{2}+\alpha\{P_{3},D\}\\ +(N\cdot H)+\eta\end{array}$ (97)
while the related determining equations (19) are reduced to the following system:
$\begin{split}&(x_{2}(\sigma x_{2}-2\nu)+\omega)M_{1}+((x_{1}(\nu-\sigma x_{2})+\kappa)M_{2}+N_{1}=0,\\ &(x_{1}(\nu-\sigma x_{2})+\kappa)M_{1}+(\sigma x_{1}^{2}+\omega)M_{2}+N_{2}=0,\\ &(\lambda^{31}-\mu x_{2}+\alpha x_{1})M_{1}+(\mu x_{1}+\lambda^{32}+\alpha x_{2})M_{2}+2\alpha M+N_{3}=0\end{split}$ (98)
where we used the shifts of variables $x_{1}$ and $x_{2}$ to nullify the maximally possible number of parameters $\lambda_{1}^{ab}$.
System (98) has non-constant and nonlinear solutions if parameters $\lambda^{3\alpha}$ are zero or can be reduced to zero by shifts. Linear solutions admit shift transformations and so have to be ignored. Thus without loss of generality these parameters can be nullified.
Equations (98) with $\lambda^{3\alpha}$ have consistent (i.e., non-invariant with respect to additional Lie groups) solutions provided we impose one of the following conditions to the remaining parameters:
$\mu\neq 0,\ \omega=0,\ \kappa=0,\ \nu=0$ (99)
$\mu=0,\ \sigma\neq 0,\ \kappa=0,\ \omega\nu=0,$ (100)
$\mu=\sigma=0,\ \nu\neq 0,\ \omega=\kappa=0,$ (101)
where we used shift transformations to nullify as many arbitrary parameters as possible.
Consider versions (99)-(101) consequently. Let arbitrary parameters satisfy conditions (99) then equations (98) are reduced to the following form:
$\begin{split}&\sigma x_{2}^{2}M_{1}-\sigma x_{1}x_{2}M_{2}+N_{1}=0,\\ &\sigma x_{1}^{2}M_{2}-\sigma x_{1}x_{2}M_{1}+N_{2}=0,\\ &(\alpha x_{1}-\mu x_{2})M_{1}+(\mu x_{1}+\alpha x_{2})M_{2}+2\alpha M+N_{3}=0\end{split}$ (102)
Solutions of system (102) for the cases $\alpha=0$ and $\alpha\neq 0$ are represented in Items 4 and 5 of Table 3 respectively.
No $f$ $V$ Integrals of motion
1 $\frac{1}{c_{1}x_{1}+c_{2}x_{2}+2c_{3}\tilde{r}^{2}+c_{4}}$ $\frac{c_{5}x_{1}+c_{6}x_{2}+2c_{7}\tilde{r}^{2}+c_{8}}{c_{1}x_{1}+c_{2}x_{2}+2c_{3}\tilde{r}^{2}+c_{4}}$ $\begin{array}[]{c}P_{1}^{2}-P_{2}^{2}\\ +\left((c_{2}x_{2}-c_{1}x_{1}+c_{3}(x_{2}^{2}-x_{1}^{2}))\cdot H\right)\\ +c_{5}x_{1}-c_{6}x_{2}+c_{7}(x_{1}^{2}-x_{2}^{2}),\\ P_{1}P_{2}-\frac{1}{2}\left(((2c_{3}+c_{2})x_{1}+c_{1}x_{2})\cdot H\right)\\ +\frac{1}{2}((2c_{7}+c_{6})x_{1}+c_{5}x_{2})\end{array}$
2 $\frac{1}{c_{1}x_{1}+F(x_{2})}$ $\frac{c_{2}x_{1}+\tilde{F}(x_{2})}{c_{1}x_{1}+F(x_{2})}$ $\begin{array}[]{c}P_{1}^{2}-P_{2}^{2}+c_{1}x_{1}-\tilde{F}(x_{2})\\ +\left((F(x_{2})-c_{1}x_{1})\cdot H\right),\\ 2P_{3}P_{1}-(c_{1}x_{3}\cdot H)+c_{2}x_{3},\\ \{P_{3},L_{2}\}-\frac{c_{1}}{2}(x_{3}^{2}\cdot H)+\frac{1}{2}c_{2}x_{3}^{2}\end{array}$
5 $\frac{\tilde{r}^{2}}{c_{1}\tilde{r}^{2}+F(\varphi)}$ $\frac{c_{2}\tilde{r}^{2}+G(\varphi)}{c_{1}\tilde{r}^{2}+F(\varphi)}$$\begin{array}[]{c}\{P_{3},D\}-2c_{1}(x_{3}\cdot H)+2c_{2}x_{3},\\ L_{3}^{2}-(F(\varphi)\cdot H)-G(\varphi),\\ D^{2}-c_{1}(r^{2}\cdot H)+c_{2}r^{2}\end{array}$
6 $\frac{\tilde{r}^{2}}{c_{1}(\mu\ln(\tilde{r})-\nu\varphi)+c_{2}\tilde{r}^{2}+c_{3}}$ $\frac{c_{4}\mu(\ln(\tilde{r})-\nu\varphi)+c_{5}\tilde{r}^{2}+c_{6}}{c_{1}\mu(\ln(\tilde{r})-\nu\varphi)+c_{2}\tilde{r}^{2}+c_{3}}$ $\begin{array}[]{c}\mu P_{3}L_{3}+\nu\{P_{3},D\}\\ -\left(2c_{2}\nu x_{3}\cdot H\right)+2c_{5}\nu x_{3},\\ L_{3}^{2}+2(c_{1}\varphi\cdot H)-2c_{4}\varphi\end{array}$
The next versions to be considered correspond to the trivial parameter $\mu$. In the case (100) equations (98) look as follows:
$\begin{split}&x_{2}(\sigma x_{2}-2\nu)M_{1}+x_{1}(\nu-\sigma x_{2})M_{2}+N_{1}=0,\\ &x_{1}(\nu-\sigma x_{2})M_{1}+\sigma x_{1}^{2}M_{2}+N_{2}=0,\\ &\alpha(x_{1}M_{1}+x_{2}M_{2})+2\alpha M+N_{3}=0\end{split}$ (103)
for $\omega=0$, and
$\begin{split}&x_{2}(\sigma x_{2}+\omega)M_{1}-\sigma x_{1}x_{2}M_{2}+N_{1}=0,\\ &\sigma(x_{1}^{2}M_{2}-x_{1}x_{2}M_{1})-\omega M_{2}+N_{2}=0,\\ &\alpha(x_{1}M_{1}+x_{2}M_{2})+2\alpha M+N_{3}=0\end{split}$ (104)
for $\nu=0$.
Solution of equations (103) with $\nu=0$ and $\alpha=0$ are presented in Item 2 of Table 2. If $\nu\neq 0$ and $\alpha=0$ we obtain the following solutions:
$M=c_{1}+\frac{F(\varphi)}{\tilde{r}^{2}},\ N=-2\nu c_{1}x_{3}-\sigma F(\varphi).$ (105)
Function $M$ (108) solves the related equations (103) with arbitrary parameters $\nu$ and $\sigma$. In particular we can set $\nu=1,\sigma=0$ or $\nu=0,\sigma=1$, and so the PDM Schrödinger equation with the mass presented in (108) admits as minimum two integrals of motion, i.e,
$\{P_{3},D\}-2c_{1}(x_{3}\cdot H)+2c_{2}x_{3}$ (106)
@@
$L_{3}^{2}-(F(\varphi)\cdot H)-G(\varphi),$ (107)
see Item 6 of Table 3.
Analogously, for $\nu\neq 0$ and $\alpha\neq 0$ we come to solutions presented in Item 7 of Table 3. The corresponding mass function is a particular case of function $M$ presented in (108).
Consider now equations (104). If parameters $\sigma,\omega$ and $\alpha$ are nontrivial we easy obtain solutions presented in Item 4 of Table 3, if $\alpha=0$ we have solutions presented in Item 3 of the same table.
The most complicated case is $\alpha=0$ and nontrivial $\sigma$ and $\omega$. Up to scaling of variables $x_{a}$ we can set $\sigma=\pm\omega=1$ and reduce (104) to the following form:
$\begin{split}&(x_{2}^{2}\pm 1)M_{1}-x_{1}x_{2}M_{2}+N_{1}=0,\\ &(x_{1}^{2}\mp 1)M_{2}-x_{1}x_{2}M_{1}+N_{2}=0,\\ &N_{3}=0\end{split}$ (108)
Excluding variable $N$ we come to the following consistency condition of this system:
$(x_{2}^{2}-x_{1}^{2}\pm 2)M_{12}+x_{1}x_{2}(M_{11}-M_{22})+3(x_{2}F_{1}-x_{1}F_{2})=0$ (109)
which looks rather cumbersome. However, choosing new dependent and independent variables, namely,
$\tilde{M}=\tilde{\omega}M,\ \tilde{x}_{1}=\tilde{\omega}+x_{1}^{2}-x_{2}^{2}\pm 1,\ \tilde{x}_{2}=\tilde{\omega}+x_{2}^{2}-x_{1}^{2}\mp 1$ (110)
where $\tilde{\omega}=\sqrt{\tilde{r}^{4}+4(x_{2}^{2}-x_{1}^{2}\mp 1)}$ we can reduce (117) to the standard D’Aalembert equation for $\tilde{M}=\tilde{M}(\tilde{x}_{1},\tilde{x}_{2})$ whose solutions are
$\tilde{M}=F_{1}(\tilde{x}_{1})+F_{2}(\tilde{x}_{2})$ (111)
where $F_{1}(.)$ and $F_{2}(.)$ are arbitrary functions. Just these solutions are presented in Item 8 of Table 3.
For any fixed functions $F_{1}(.)$ and $F_{2}(.)$ we can solve equation (108) and find the corresponding function $N$. Unfortunately it is impossible to find $N$ in closed form for $F_{1}(.)$ and $F_{2}(.)$ arbitrary. However, it can be done for the special class of these functions presented in Item 8 of Table 2.
Let us represent the system of determining equations (98) corresponding to the remaining version (101):
$\begin{array}[]{l}\nu(x_{1}M_{2}-2x_{2}M_{1})+N_{1}=0,\\ \nu x_{1}M_{1}+N_{2}=0,\\ \alpha(x_{1}M_{1}+x_{2}M_{2}+2M)+N_{3}=0\end{array}$ (115)
Their solutions for $\alpha=0,\nu\neq 0$, $\alpha\neq 0,\nu=0$ and $\alpha\nu\neq 0$ are represented in Item 3 of Table 3, Item 6 of Table 4 and Item 3 of Table 5 respectively.
5.3 Systematic search for superintegrable systems
In the previous section we searched for integrable systems which admit one second order integral of motion. Just such systems are collected in Table 3.
However some of the found systems accidentally admit two independent second order integrals of motion and so are superintegrable. Such systems are collected in Table 4, see Items 2-7 there.
But it is necessary to provide a systematic search for superintegrable and maximally superintegrable systems. To do it we will start with the solutions found in Sections 5.1 and 5.2, and test their compatibility with equations (98) for all admissible sets of arbitrary parameters $\alpha,\mu,\nu,\sigma,\omega,\kappa,\lambda^{31},\lambda^{32}$ discussed in the above.
Let us start with the system whose arbitrary parameters are represented in Item 1 of Table 3, refer to equations (88) and (89). They satisfy equations (87) with $\lambda^{31}=\lambda^{32}=\nu=0$ and $\mu=1$. Our goal is to verify if they can satisfy these equations with another values of these parameters.
We already know that it is the case if the generic potential (89) is reduced to the forms presented in (91) and (93). The related equations (87) include nontrivial parameters $\lambda^{3\alpha}$ while parameter $\nu$ was proposed to be trivial. Notice that the triviality of this parameter is supported by the possibility to reduce it to zero using the rotation transformation.
But what has happen wether the system admits a pair of integrals of motion of generic form (85)? In general we can nullify the coefficient $\nu$ only in one of them, while the second integral can include the term $\nu P_{1}P_{2}$. In other words, the system admitting integral of motion (90) and so including the mass function $M$ presented in (88) can admit one more integral of motion(85) provided this function satisfies equations (87) with $\nu\neq 0$ but $\mu=0$. Solving these equations for all possible values of arbitrary parameters we come to the following versions.
$\lambda^{31}=\lambda^{32}=0,\ M=c_{1}x_{1}+c_{2}x_{2}+c_{3}\tilde{r}^{2}+c_{4},\ V=c_{5}x_{1}+c_{6}x_{2}+c_{7}\tilde{r}^{2}+c_{8},$ (116)
$(\lambda^{31})^{2}+(\lambda^{32})^{2}\neq 0,\ M=c_{1}x_{1}+c_{2}x_{2}+c_{3},\ V=c_{4}x_{1}+c_{5}x_{2}+c_{6}.$ (117)
Version (117) had been already discussed in Section 5.1. Integrals of motion for the PDM systems with arbitrary elements (116) are represented in Item 2 of Table 4.
The next step is to verify wether the PDM system with arbitrary elements (88) and (89) can admit additional integral of motion of generic form (97). To do it we are supposed to solve the determining equations (98) for function $M$ represented in (88). Substituting (88) into (98) we come to the following system:
$\begin{split}&(\sigma x_{2}^{2}-\nu x_{2})F_{1}+(\nu x_{1}+\kappa-\sigma x_{1}x_{2})G_{2}+N_{1}=0,\\ &\sigma x_{1}^{2}G_{2}+((\nu-\sigma x_{2})x_{1}+\kappa)F_{1}+N_{2}=0,\\ &(\alpha x_{1}-\mu x_{2})F_{1}+(\mu x_{1}+\alpha x_{2})G_{2}+N_{3}=0.\end{split}$ (118)
Notice that whenever parameter $\sigma$ is not trivial it is possible to reduce parameter $\nu$ to zero making a shift of variable $x_{2}$, so we can impose the condition $\sigma\nu=0$.
Equations (118) have nonconstant solutions for $M$ provided their parameters satisfy one of the following conditions:
$\nu=0,\alpha=0,\mu\neq 0,\sigma\neq 0$ (119)
$\nu=0,\alpha=0,\mu=0,\kappa=0,\sigma\neq 0,$ (120)
$\nu=0,\alpha\neq 0,\mu=0,\kappa=0,\sigma\neq 0,$ (121)
$\sigma=0,\mu=0,\kappa=0,\alpha=0,\nu\neq 0,$ (122)
$\sigma=0,\mu=0,\kappa=0,\alpha\neq 0,\nu\neq 0$ (123)

### v3

@@ 4 PDM systems invariant with respect to shifts
Let Q be an integral of motion (4) admitted by equation (1) with arbitrary elements (6). By definition $P_{3}$ is the integral of motion too, the same is true for the commutators $[P_{3},Q],[P_{3},[P_{3},Q]]$ and $[P_{3},[P_{3},[P_{3},Q]]].$ Thus any second order symmetry induces the symmetry generated by Killing tensors $\mu^{ab}$ (12)-(20) independent on $x_{3}$ which are itemized below together with one additional special tensor:
@@
Classifying PDM systems admitting these induced symmetries we can obtain the complete list of the systems invariant with respect to shifts and admitting second order integrals of motion. Then it will be necessary to specify the additional integrals of motion which can be admitted by the found systems. As a result we obtain the complete list of the systems presented in the following Tables 2-4. The detailed calculations are presented in Sections 4.1 - 4.3.
4.1 Integrals of motion whose differential part $Q_{(n)}$ (22) is independent on $x_{3}$
@@
The first step of our programm is to classify the systems whose integrals of motion are generated by the Killing tensors specified at the beginning of Section 4, i.e., by the following linear combinations:
@@
$Q=Q^{(1)}+Q^{(2)}$ (95)
where
$Q^{(1)}=\omega P_{1}^{2}+\kappa P_{1}P_{2}+\nu\{P_{1},L_{3}\}+\sigma L_{3}^{2}+(N^{(1)}\cdot H)+\eta^{(1)},$ (96)
$Q^{(2)}=\lambda^{3\alpha}P_{3}P_{\alpha}+\mu L_{3}P_{3}+\alpha\{P_{3},D\}+(N^{(2)}\cdot H)+\eta^{(2)}.$ (97)
The related determining equations (24) are reduced to the following system:
$\begin{split}&(x_{2}(\sigma x_{2}-2\nu)+\omega)M_{1}+((x_{1}(\nu-\sigma x_{2})+\kappa)M_{2}+N_{1}=0,\\ &(x_{1}(\nu-\sigma x_{2})+\kappa)M_{1}+(\sigma x_{1}^{2}-\omega)M_{2}+N_{2}=0,\end{split}$ (98)
$(\lambda^{31}-\mu x_{2}+\alpha x_{1})M_{1}+(\mu x_{1}+\lambda^{32}+\alpha x_{2})M_{2}+2\alpha M+N_{3}=0.$ (99)
Equations (98), (99) include eight arbitrary parameters. But in fact their number can be essentially reduced using shift transformations and analysing this system consistency. Just such reduction is the subject of the following subsection.
The specific property of system (98), (99) is that that equations (98) include the set of arbitrary parameters $\sigma,\nu,\omega,\kappa$ and do not include $\alpha,\mu,\lambda^{31},\lambda^{32}$ while equation (99) includes the second set of parameters and do not include the first one. It means that system (98), (99) with $\sigma=\nu=\omega=\kappa=0$ represents the determining equations for integral of motion $Q^{(2)}$. Moreover, in the case when $\alpha,\mu,\lambda^{31}$ and $\lambda^{32}$ are trivial but some of the other coefficients are nonzero we have the determining equations for $Q^{(1)}$ In the another combinations of the arbitrary parameters the discussed system defines the PDM Schrödinger equations admissing both integrals of motion, i.e., $Q^{(1)}$ and $Q^{(2)}$.
4.1.1 Systems admitting integral of motion $Q^{(2)}$
Let us start with the case when parameters $\sigma,\nu,\omega$ and $\kappa$ are trivial. The corresponding subsystem (98) is reduced to the conditions $N_{1}=N_{2}=0$ so the related function $N$ can depend on $x_{3}$ only. Moreover, since the l.h.s. of equations (99), does not depend on $x_{3}$, the generic form of function $N$ looks as follows:
$N=2\lambda x_{3}+G(x_{1},x_{2})$ (100)
where $\lambda$ is a constant.
Substituting (100) into (99) we come to the following equation:
$(\lambda^{31}-\mu x_{2}+\alpha x_{1})M_{1}+(\mu x_{1}+\lambda^{32}+\alpha x_{2})M_{2}+2\alpha M+2\lambda=0.$ (101)
Let at least one of parameters $\mu$ or $\alpha$ be nonzero. Then, using shifts of variables $x_{1}$ and $x_{2}$ we can reduce to zero $\lambda^{31}$ and $\lambda^{32}$. After that we can integrate equation (101) and obtain the generic form of the corresponding function $M$ :
$M=\frac{F(\alpha\ln(\tilde{r})-\mu\varphi)}{\tilde{r}^{2}}-\frac{\lambda}{\alpha},\ \alpha\lambda\neq 0$ (102)
$M=F(\tilde{r})-2\lambda\varphi,\ \alpha=0,\ \mu\neq 0,$ (103)
@@ 4.1.1 Systems admitting integral of motion $Q^{(2)}$
$M=\frac{F(\varphi)}{\tilde{r}^{2}}-\frac{\lambda}{\alpha},\alpha\neq 0,\ \mu=0.$ (104)
For both $\alpha$ and $\mu$ trivial we cannot a priori reduce reduce $\lambda^{31}$ and $\lambda^{32}$ to zero, but can normalize them by simultaneous scaling of all independent variables $x_{1},x_{2}$ and $x_{3}$. The corresponding equation (101) again can be solved exactly. The related generic form of $M$ is:
$M=-\lambda y_{1}+F(y_{2})$ (105)
where $y_{1}=\lambda^{31}x_{1}+\lambda^{32}x_{2},\ y_{2}=\lambda^{31}x_{2}-\lambda^{32}x_{1}$ and the related parameters $\alpha,\mu,\lambda^{31}$ and $\lambda^{32}$ are supposed to satisfy the following condition:
$\alpha=0,\ \mu=0,(\lambda^{31})^{2}+(\lambda^{31})^{2}=1.$ (106)
Using rotations around the third coordinate axis we can reduce $y_{1}$ to $x_{1}$ and $y_{2}$ to zero.
Thus we fix four inequivalent versions of parameters $\alpha,\mu,\lambda^{31}$ and $\lambda^{32}$ and find the generic form of the corresponding mass functions $M$ which are presented in equations (102), (103) (104) and (106). The corresponding potentials are easy calculated using definition (25).
The related PDM system admit second order integrals of motion presented in (97) where $N=c_{1}x_{3}$. In other words, these integrals of motion have the following forms:
$Q^{(2)}=\alpha\{P_{3},D\}+\mu L_{3}P_{3}+2\lambda(x_{3},\cdot H)+\lambda^{31}P_{3}P_{1}-2\tilde{\lambda}x_{3}$ (107)
where parameters $\alpha$ and $\mu$ take the values fixed in (102) - (104).
4.1.2 Systems admitting integral of motion $Q^{(1)}$
One more possibility corresponds to the case when all parameters $\alpha,\mu,\lambda^{31}$ and $\lambda^{32}$ are trivial, and system (98), (99) is reduced to its subsystem (98) added by the condition $N_{3}=0.$ As in the previous subsection we have four arbitrary parameters only. Moreover, their number of this parameters can be reduced. Indeed, let parameter $\sigma$ is nontrivial making the shift $x_{2}\to x_{2}+\frac{\nu}{\sigma}$ we can nullify parameter $\nu$. If $\sigma$ is trivial but $\nu$ is nonzero we can nullify parameter $\omega$ by the shift $x_{2}\to x_{2}-\frac{\omega}{2\nu}$.
In all the cases discussed above it is also possible to nullify parameter $\kappa$ making the shift $x_{1}\to x_{1}-\frac{\kappa}{\nu}.$ Whenever both $\sigma$ and $\nu$ we also can nullify $\kappa$ making the rotation around the third coordinate axis.
In other words, there are the following inequivalent versions of parameters $\sigma,\nu$ and $\omega$ which have to be considered:
$\alpha=\mu=\lambda^{31}=\lambda^{32}=\kappa=0.$ (108)
Moreover, conditions (108) are added by one of the following ones:
$\sigma=0,\nu\neq 0,\ \omega=0,$ (109)
$\sigma\neq 0,\nu=0,\ \omega=0,$ (110)
$\sigma=0,\nu=0,\ \omega\neq 0$ (111)